OpenStudy (anonymous):

whats the P(X1+X2=1)?

OpenStudy (anonymous):



OpenStudy (zarkon):

\[P(X_1+X_2=1)\] \[=P([X_1=0,X_2=1] \text{ or } [X_1=1,X_2=0])\]

OpenStudy (anonymous):

can you explain more?

OpenStudy (anonymous):

In which cases can X1+X2 amount to one? Look at the table and see that it happens only when: (X1=1 AND X2=0) OR (X1=0 AND X2=1) P [(X1=1 AND X2=0) OR (X1=0 AND X2=1)]=P(X1=1 AND X2=0)+P(X1=0 AND X2=1)= 2/12+3/12=5/12